OpenStudy (anonymous):

What is -5/8 divide by 9/10 ? A. -25/36 Dividing Fractions B. -9/16 Dividing Fractions C. -45/80 Dividing Fractions D. -1/36 Dividing Fractions

OpenStudy (anonymous):

it is c

OpenStudy (anonymous):

i am just making sure if i am right

OpenStudy (anonymous):

\[\frac{ -5 }{ 8 }\div \frac{ 9 }{ 10 }\] So since we are dividing by fractions you want to multiply by the reciprocal so it would become \[\frac{ -5 }{ 8 }\times \frac{ 10 }{ 9 }\] which would then multiply to become \[\frac{ -50 }{ 72 }\] This value can be simplified to\[\frac{ -25 }{ 36 }\] which would make the answer A

OpenStudy (charlotte123):

@studyislandhelp it is A :D -5/8 / 9/10 = -25/36 ^_^

OpenStudy (anonymous):

The answer would be A You have to set it up as follows -(5/8) divided by (9/10) -.625 divided by .9 you get -.694 and -25/36 equals -.694 do you understand?